Assume the bid rate of a Singapore dollar is $.40 while the ask rate is $.41 at Cold Bank. Assume the bid rate of a Singapore dollar is $.42 while the ask rate is $.425 at Red Dot Bank. Given this information, what would be your gain if you use $1,575,000 and execute loca­tional arbitrage? That is, how much will you end up with over and above the $1,575,000 you started with?
Answer #1

Rate at Cold Bank 1S$ = $0.40 - $0.41

Rate at Red dot Bank 1S$ = $0.42 - $0.425

Purchase S$ from Cold Bank @$0.41 for $ 1,575,000

S$ from Cold Bank = 1575000 / 0.41

= 3,841,463.41

Sell These S$ in Red dot bank.

= 3,841,463.41 * 0.42

= $ 1,613,414.63

Book Profit = 1613414.63 - 1575000

= $ 38,414.63

Pls comment, if any further assistance is required